OpenStudy (hulkart):

Help....;) Event A has a 0.3 probability of occuring and event B has a 0.4 probability of occuring. A and B are independent events. What is the probability that neither A or B occurs? 0.42 0.10 0.70 0.12

OpenStudy (hulkart):

so....whats the answer from the choice?

OpenStudy (hulkart):

so, 0.70?

OpenStudy (hulkart):

haha yes of course : /

OpenStudy (kropot72):

\[\large P(\bar {A})=1-0.3=0.7\] \[\large P(\bar{B})=1-0.4=0.6\] \[\large P(\bar{A} \cap \bar{B})=P(\bar{A}) \times P(\bar{B})=0.7\times0.6=you\ can\ calculate\]

OpenStudy (hulkart):

you're the real mvp +kropot72 so, it's 0.42

OpenStudy (kropot72):

Yes, 0.42 is correct.

OpenStudy (kropot72):

The required probability is the intersection of the probabilities of A not occurring and B not occurring.